OpenStudy (anonymous):

find an nth-degree polynomial function with real coefficients satisfying the given conditions. If you are using a graphing utility, use it to graph the function and verify the real zeros and the given function value n=3;4 and 4 i are zeros; f (-1)=-170

OpenStudy (cruffo):

Sounds like a system of equations problem. We are looking for a 3rd degree polynomial because n = 3. So let \[f(x) = ax^3 + bx^2 + cx + d.\] Since there are four unknowns, we need four equations in order to solve this.

OpenStudy (cruffo):

Since 4 is a root, you have that \[f(4) = 0.\] Thus equation 1 is \[64a + 16b + 4c + d = 0\]. Also, you have that \[f(-1) = -170.\] So equation 2 is \[-a+b-c+d = -170.\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would I put the polynomial in standard form

OpenStudy (cruffo):

Wait, you still need two more equations!!

OpenStudy (cruffo):

Since 4i is a root, that automatically means -4i is also a root (complex roots always travel in pairs) So the factored form would look like \[(x+4i)(x-4i).\] When multiplied out this gives the polynomial \[x^2 +16.\] Hum...

OpenStudy (cruffo):

Maybe we don't need the system stuff... If we think of the polynomial in factored form \[f(x) = a(x+r_1)(x+r_2)(x+r_3)\] Use the fact that that 4 , 4i, and -4i are roots as above, \[f(x) = a(x-4)(x^2+16)\] Then we would be able to use the fact that \[f(-1) = a(-5)(17) = -170\] to solve for a.

OpenStudy (anonymous):

ok

OpenStudy (cruffo):

so a = 2. \[f(x) = 2(x-4)(x^2+16)\] Now all that's left to do is multiply the factors, which will give you the standard form.

OpenStudy (anonymous):

so it 2x-8+x^2+16

OpenStudy (cruffo):

almost, looks like you typed addition instead of multiplication. \[(2x-8)(x^2+16)\]

OpenStudy (anonymous):

ya my so it in standard form now rite

OpenStudy (anonymous):

i mean my fault

OpenStudy (cruffo):

No, you still need to multiply the factors. The goal is to have a polynomial with no parenthesis. Do you remember "FOIL"?

OpenStudy (anonymous):

yes

OpenStudy (cruffo):

That's what you need! \[(2x-8)(x^2+16)\] First, outer, inner, last \[2x^3 + 32x - 8x^2 - 128\] Now write in descending order \[2x^3 -8x^2 + 32x - 128\] To check that the answer is correct, find f(-1) as it should be -170. done :)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

hello